Circular No. 6110

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) Phone 617-495-7244/7440/7444 TWX 710-320-6842 ASTROGRAM CAM

MARSDEN@CFA.HARVARD.EDU or GREEN@CFA.HARVARD.EDU (science)

SUPERNOVA 1994af IN ANONYMOUS GALAXY R. H. McNaught, Anglo-Australian Observatory, reports his discovery

of an apparent supernova (mag ∼ 18) at α = 3^{h}02^{m}57^{s}.97, δ = −12^{o}22′54′′.9 (equinox 1950.0, uncertainty 0′′.4 in each coordinate), or ∼ 3′′ due east of the center of the host spiral galaxy. SN 1994af was found on an R plate taken on Nov. 27 by C. P. Cass with the U.K. Schmidt Telescope. No image appears at this position on another R plate taken on 1993 Sept. 23 or on the SERC J survey. A nearby star (∼ 0.5 mag fainter than SN 1994af) has position end figures 3^{m}05^{s}.52, 8′′.6.

B. Schmidt, P. Challis, and R. Kirshner, Harvard-Smithsonian Center for Astrophysics (CfA), report that a spectrogram (380–740 nm) of SN 1994af taken by P. Berlind (CfA) with the Tillinghast 1.5-m reflector (+ FAST spectrograph) on Dec. 1.25 UT shows this object to be a relatively young type-II supernova, as indicated by broad P-Cyg profiles of Hα and He I (587.6 nm). The velocity of the host galaxy, measured from narrow Hα emission, is 3750 km/s.

SUPERNOVA 1994ae IN NGC 3370 S. Nakano, Sumoto, Japan, forwards the following accurate position

from Y. Kushida and R. Kushida at Yatsugatake South Base Observatory (from a CCD image taken on Nov. 30.842 UT, when the supernova was at unfiltered mag 12.5): α = 10^{h}44^{m}21^{s}.52, δ = +17^{o}32′20′′.7 (equinox 1950.0).

OJ 287 H. Inoue, F. Nagase, F. Makino, C. Otani, R. Fujimoto, and Y. Ueda,

Institute of Space and Astronautical Science; and M. Tashiro, T. Takahashi, and E. Idesawa, University of Tokyo, communicate on behalf of the ASCA Team: “We observed OJ 287 with ASCA between Nov. 18.42 and 19.62 UT, with an exposure time of ∼ 40 000 s, and detected significant x-ray emission from the source. The flux was then (4.2 ± 0.2) × 10^{−12} erg cm^{−2} s^{−1} in the range 2–10 keV, which is smaller by about a factor of two than when seen with the Einstein Observatory in 1979 and 1980 (Madjeski and Schwartz 1988, Ap.J. 330, 776), in spite of the current high luminosity at optical bands (IAUC 6095). The intensity variation during our observation was < 30 percent on time scales from 100 min to a day. The 0.5- to 10-keV spectrum is fit by a power-law spectrum with photon index 1.7 ± 0.1 and with interstellar-absorption column density < 1 × 10^{21} cm^{−2}.”

1994 December 3 Daniel W. E. Green